OpenStudy (anonymous):

1. What is the maximum height that a basketball will have if thrown up into the air from a height of six feet, with an initial speed of 20 feet per second? (Must use quadratic formula). 2. After how many seconds does the basketball hit the ground? (Must solve by completing the square)

rvc (rvc):

Initial velocity,u=20m/s

rvc (rvc):

considering upwards positive then acceleration due to gravity is negative

rvc (rvc):



rvc (rvc):

considering v=u+at since at h_max v=0 that is final velocity will be 0 therefore u will get time... as u=20feet/sec g=9.81 m/s there has to be a conversion ....

rvc (rvc):

or else v^2=u^2+2 a s

rvc (rvc):

well @Joe-Palooka do you have any idea?

rvc (rvc):

like u might have solved similar question

rvc (rvc):

its a simple physics kinematics problem :/

OpenStudy (anonymous):

Hi. Thanks, but I am still lost, since I need to use the quadratic equation on the first part and then use "completing the square" for part two. I think I have part one done.

rvc (rvc):

please show us the working for part 1

rvc (rvc):

@Astrophysics

OpenStudy (astrophysics):

\[v_0 = 20 ft/s,~t=0,~h_0= 6 ft\] \[y(t) = -g t^2+v_0t+h_0\] this is your equation for a free falling body

OpenStudy (astrophysics):

Or you may use \[y(t) = -16t^2+v_0t+h_0\] where that g is in ft/s^2

OpenStudy (astrophysics):

To find the max height you can just use the vertex, but understand what it is \[h = \frac{ -b }{ 2a }\]

OpenStudy (anonymous):

Thank you all. Very helpful.

rvc (rvc):

my pleasure... All the best :)

OpenStudy (anonymous):

g=-32 ft/sec u=20 ft/s v=0 ft /s \[v=u+\gt,0=20-32t,t=\frac{ 20 }{ 32 }=\frac{ 5 }{ 8 } \sec,h=ut+\frac{ 1 }{ 2 }g t^2+6\] \[h=20 \times \frac{ 5 }{ 8 }+\frac{ 1 }{ 2 }\times \left( -32 \right)\left( \frac{ 5 }{ 8 } \right)^2+6\] =?

OpenStudy (anonymous):

correction v=u+ gt not v=u+>

OpenStudy (mrnood):

the direct solution is obtained from the equation v^2 = u^2 +2as v is velocity at peak (=0) u is intitial velocity = 20 ft/s a = g = -32ft/s^2 s is your solution - then add the intial 6 ft

OpenStudy (cheesecakekitten):

http://openstudy.com/study#/updates/5651f61de4b0aa790266f390

jagr2713 (jagr2713):

Great job guys ! Thanks for helping